OpenStudy (anonymous):

- 1/2x = - 9/10 solve using the multiplication principle and simplify your answer.

OpenStudy (asadkarim7):

cross multilpy 10 (-1x) = 2 (-9) -10x = - 18 x= -18/-10 x= 1.8

OpenStudy (anonymous):

are you supposed to flip the equations?

OpenStudy (asadkarim7):

yeah

OpenStudy (asadkarim7):

what else can we do

OpenStudy (anonymous):

I didn't know on these ones you are supposed to do that. Thank you.

OpenStudy (anonymous):

what do you mean what else can we do ?

OpenStudy (asadkarim7):

i mean there is nothing else we can do to it :P

jhonyy9 (jhonyy9):

-10= -18x --- x=10/18 = 5/9